OpenStudy (anonymous):

28 students is 70% of how many students?

OpenStudy (baldymcgee6):

x*0.70 = 28

OpenStudy (anonymous):

How did you get 28

OpenStudy (baldymcgee6):

It was given... 70% of x is 28

OpenStudy (anonymous):

so the correct answer is 28 students

OpenStudy (anonymous):

or do i do 0.70 * 28

OpenStudy (baldymcgee6):

28 students was given to us in the question... solve for x in the equation i gave you

OpenStudy (anonymous):

i did 28 * 0.70= 19.6

OpenStudy (baldymcgee6):

\[x*0.70 = 28\] Divide both sides by 0.70 to solve for x \[x* \frac{ 0.70 }{ 0.70 } = \frac{ 28 }{ 0.70 }\] \[x = \frac{ 28 }{ 0.70 } = 40 students\]

OpenStudy (anonymous):

oh ok. so i divide 28/0.70= 40

OpenStudy (anonymous):

ok. can you help me with another math problem please